# Parameter estimation in hyperbolic linear SPDEs from multiple measurements

### v1

@@ Abstract
The coefficients of elastic and dissipative operators in a linear hyperbolic SPDE are jointly estimated using multiple spatially localised measurements. As the resolution level of the observations tends to zero, we establish the asymptotic normality of an augmented maximum likelihood estimator. The rate of convergence for the dissipative coefficients matches rates in related parabolic problems, whereas the rate for the elastic parameters also depends on the magnitude of the damping. The analysis of the observed Fisher information matrix relies upon the asymptotic behaviour of rescaled $M,N$-functions generalising the operator sine and cosine families appearing in the undamped wave equation. In contrast to the energetically stable undamped wave equation, the $M,N$-functions emerging within the covariance structure of the local measurements have additional smoothing properties similar to the heat kernel, and their asymptotic behaviour is analysed using functional calculus.
@@ 1 Introduction
for locations $x_{1},\dots,x_{N}\in\Lambda$ and $\gamma\in\{\alpha_{i},\beta_{j}|1\leq i\leq p;1\leq j\leq q\}$. The point spread functions [8, 9] $K_{\delta,x_{k}}$ are compactly supported functions taking non-zero values in an area centred around $x_{k}$ with radius $\delta$. Local measurements emerge naturally as they describe the physical limitation of measuring $u(t,x_{k})$, which, in general, is only possible up to a convolution with a point spread function. Local observations were introduced to the field of statistics for SPDEs in [5], where the authors investigated a stochastic heat equation with a spatially varying diffusivity. It was shown that the diffusivity at location $x_{k}\in\Lambda$ can be estimated based on a single local measurement process at $x_{k}$ as the resolution level $\delta$ tends to zero. The local observation scheme turned out to be robust under semilinearities [4, 3], multiplicative noise [19], discontinuities [33] or lower-order perturbation terms [6, 38]. In contrast to the estimation of the diffusivity, the identifiability of transport or reaction coefficients necessarily requires an increasing amount $N\rightarrow\infty$ of measurements. In the recent contribution [41], the local measurement approach was extended to hyperbolic problems and the non-parametric wave speed in the undamped stochastic wave equation was estimated by relating the observed Fisher information to the energetic behaviour of an associated deterministic wave equation.
@@
with $N=N(\delta)$ measurements. The consistent estimation for $\vartheta_{i}$ holds if $N^{1/2}\delta^{-2\alpha_{i}+\alpha_{1}+\beta_{1}}\rightarrow\infty$, whereas $\eta_{j}$ can be estimated in the asymptotic regime $N^{1/2}\delta^{-2\beta_{j}+\beta_{1}}\rightarrow\infty$. In particular, estimating elastic coefficients is more difficult under higher dissipation, while damping coefficients are unaffected by the order of the elastic operator $A_{\vartheta}$ and their convergence rates reflect the rates obtained in advection-diffusion equations, cf. [6]. For the maximal number of non-overlapping observations $N\asymp\delta^{-d}$, our convergence rates match the rates obtained in the spectral approach up to specific boundary cases. In the weakly damped case, i.e. $\beta_{1}=0,$ we confirm the results in [25]. That is, the dependence of the time horizon $T$ of the asymptotic variance resembles the explosive, stable and ergodic cases of the maximum likelihood drift estimator for an Ornstein-Uhlenbeck process, cf. [22, Proposition 3.46]. In the structural damped case $(\beta_{1}>0)$, the asymptotic variance is of order $T^{-1}$ instead.
@@ 2.2 The model
    Figure 2.1 displays a heatmap illustrating both the weakly and structurally damped plate equation in one spatial dimension. The solution of the SPDEs were approximated on a fine time-space grid using the finite difference scheme associated with the semi-implicit Euler-Maruyama method, see [26, Chapter 10]. Additional smoothing properties in the structurally damped case due to the dissipative operator $B_{\eta}$ result in an accelerated energetic decay in comparison to the weakly damped case.
@@ 3 The estimator
$\begin{pmatrix}\hat{\vartheta}_{\delta}\\ \hat{\eta}_{\delta}\end{pmatrix}=\begin{pmatrix}{\vartheta}_{\delta}\\ {\eta}_{\delta}\end{pmatrix}+\lVert K\rVert_{L^{2}(\operatorname{{\mathbb{R}}}^{d})}\mathcal{I}_{\delta}^{-1}\mathcal{M}_{\delta}$ (3.3)
@@
If the equation is weakly damped, i.e. if $\beta_{1}=0$ and $\eta_{1}<0$, then the term $-T\eta_{1}$ dominates the expression $(e^{T\eta_{1}}-T\eta_{1}-1)/2\eta_{1}^{2}$ in the asymptotic variance within (3.6) as $T\rightarrow\infty$. The converse is true in the amplified case with $\eta_{1}>0$. If $\eta_{1}=0$, the asymptotic variance of the augmented MLE for $\vartheta$ and $\eta_{1}$ depends on the time horizon through $T^{-2}$ as discussed in [41, Remark 5.8]. It mirrors the rate of convergence of the MLE in the ergodic, stable, and explosive case of the standard Ornstein-Uhlenbeck process as described in [22, Proposition 3.46].
@@
    Thus, the augmented MLE attains the convergence rate known from the spectral approach, see [25, 24] if it is provided with the maximal number of spatial observations $N\asymp\delta^{-d}$. Interestingly, the limiting variance of $\hat{\eta}_{\delta}$ is independent of the kernel function $K$ similar to the augmented MLE for the first order transport coefficient in [6].
@@
    holds under the structural damping given in (3.8). The asymptotic variances between $\hat{\vartheta}_{\delta}$ and $\hat{\eta}_{\delta}$ coincide in the cases $\vartheta_{1}=\lVert\Delta_{0}K\rVert^{2}_{L^{2}(\operatorname{{\mathbb{R}}}^{d})}\lVert K\rVert_{L^{2}(\operatorname{{\mathbb{R}}}^{d})}^{-2}$ or $\vartheta_{1}=1$, respectively. The consistency and the varying convergence rates of the estimators are visualised in Figure 3.1. Based on the finite difference scheme within the semi-implicit Euler-Maruyama method [26, Chapter 10] (with $10000000\times 2000$ time-space grid points), we computed the root mean squared error (RMSE) for decreasing resolution level $\delta$ from 100 Monte Carlo runs, $N\asymp\delta^{-1}$ measurement locations and the kernel function $K(x)=\exp(-5/(1-x^{2}))\mathbf{1}(|x|<1)$. In the weakly damped case, it can be seen that the estimator for the elastic coefficient $\vartheta_{1}$ achieves a much quicker convergence rate than the estimator of the damping coefficient $\eta_{1}$. On the other hand, their rates are equal under structural damping. The asymptotic variances are attained in both cases.
@@
    hold. Similar results were found in the spectral regime, cf. [24, Theorem 1.1 and Theorem 4.1]. Interestingly, the authors verified a slightly stronger consistency condition, resulting in a logarithmic rate under equality in (3.10) and (3.11). Otherwise, they also obtain the rates in (3.9). We believe that the logarithmic rates in the boundary cases are also valid in the local measurement approach given that less restrictive assumptions on the kernel $K$ are imposed, similar to [6, Proposition 5.4] in a related parabolic problem.
@@ 4 Proofs
We will frequently use that $L_{\vartheta,\eta}$ is an (unbounded) normal operator with spectrum $\sigma(L_{\vartheta,\eta})$ and the resolution of identity $E$ (cf. [34, Chapter 13]). By the functional calculus for normal unbounded operators, we can define the operator $f(L_{\vartheta,\eta})\coloneqq\int_{\sigma(L_{\vartheta,\eta})}f(\lambda)\mathop{}\!\mathrm{d}E(\lambda)$ on the domain
@@
Suppose $z\in\dot{H}^{2\alpha_{1}}(\Lambda_{\delta,x})$ such that $z_{\delta,x}\in\dot{H}^{2\alpha_{1}}(\Lambda)\subset\mathcal{D}_{f}$. Then, the claim follows immediately for $f(x)=x$ by differentiating $z_{\delta,x}$ and from the definition of $L_{\vartheta,\eta,\delta,x}$, see also [4, Lemma 16]. Using (i) and (iii) of [35, Proposition 5.15], the result can be extended to measurable $f:\mathbb{C}\rightarrow\mathbb{C}\cup\{\pm\infty\}$ by first passing to the associated resolution of the identities of $L_{\vartheta,\eta},B_{\eta},A_{\vartheta}$ and $L_{\vartheta,\eta,\delta,x},B_{\eta,\delta,x},A_{\vartheta,\delta,x}$ respectively, and interpreting the localisation as a bounded linear operator from $L^{2}(\Lambda_{\delta,x})$ to $L^{2}(\Lambda)$. ∎
@@
If $L_{\vartheta,\eta}$ is not a positive operator and has non-positive eigenvalues, any choice of the operator root is a complex-valued operator. Consequently, the associated family of $M,N$-functions in the following subsection is again complex-valued. Thus, inner products hereinafter are associated with complex Hilbert spaces. However, this does not influence the asymptotic results of Section 4.2 due to the convergence $L_{\vartheta,\eta,\delta,x}\rightarrow\bar{L}_{\vartheta,\eta}$ to a positive limiting operator $\bar{L}_{\vartheta,\eta}$.
@@ 4.1 Properties of generalised cosine and sine operator functions
Note that by [35, Theorem 5.9] all of the appearing operators $e^{B_{\eta}t/2}$, $\cos(tL_{\vartheta,\eta}^{1/2})$, $\sin(tL_{\vartheta,\eta}^{1/2})$, $B_{\eta}$ and $L_{\vartheta,\eta}^{-1/2}$ in (4.2) and (4.3) are well-defined and even commute on the smallest occurring domain as they are all based on the same underlying Laplace operator. By direct computation, one can now verify that the conditions (M1)-(M4) in [29, Definition 1.7.2] are satisfied by $M(t)$ and $N(t)$ from (4.2) and (4.3) using the functional calculus. ∎
@@
The unique positive self-adjoint operator root of the positive self-adjoint operator $L_{\vartheta,\eta}$ is well-defined and exists by [35, Proposition 5.13]. Thus, in view of Lemma 4.4, the $M,N$-functions can each be interpreted as the applications of a real-valued function to the underlying Laplace operator on a bounded spatial domain. In particular, by [35, Theorem 5.9] the $M,N$-functions are self-adjoint. ∎
@@
Let $\beta_{1}>0.$ We only prove (4.7), since the argument for (4.6) is similar, using additionally that $L_{\vartheta,\eta,\delta,x}$ commutes with $(-\Delta_{\delta,x})^{\gamma}$, $\operatorname{ord}(L_{\vartheta,\eta,\delta,x})=2\alpha_{1}$ and a bound of $L_{\vartheta,\eta,\delta,x}$ in terms of its leading term $(-\Delta_{\delta,x})^{\alpha_{1}}.$ Let $(e_{k})_{k\in\operatorname{{\mathbb{N}}}}$ form an orthonormal basis of $(-\Delta)$ in $L^{2}(\Lambda)$ with eigenvalues $\lambda_{k}>0$. Then, there exists a constant $c(\Lambda)$ such that $\lambda_{k}\geq c(\Lambda)$ for all $k\geq 1$, see [36, Proposition 5.2 and Corollary 5.3]. We consider the most involved case, that is, $\eta_{1}<0$ and $\eta_{2},\dots,\eta_{q}>0$. Consequently, $B_{\eta}$ will, in general, not be a negative operator, but there exists $y_{0}>0$ such that for all $y\geq y_{0}$, we have
@@
Thus, all eigenvalues of the operator difference $B_{\eta,\delta,x}-\delta^{2\beta_{1}}C_{1}$ are negative and the difference is bounded by $\eta_{1}(-\Delta_{\delta,x})^{\beta_{1}}/2$ in the sense that
@@
The theory of cosine operator functions was developed by Sova, [37] and led to general deterministic solution theory for undamped second-order abstract Cauchy problems. By substituting the time derivative as its own variable, it is possible to rewrite a second-order abstract Cauchy problem as a first-order abstract Cauchy problem in two components. The associated strongly continuous semigroup then lives on a product of Hilbert spaces called the phase-space; see [7, Chapter 3.14], [30] and [29, Chapter 0.3]. The same remains true under suitable assumptions on the elastic and dissipation operator within a damped abstract second-order Cauchy problem [29, Chapter 1.7].
@@
It is well-known that in the special case where both $A_{\vartheta}$ and $B_{\eta}$ are strictly negative operators $\mathcal{A}_{\vartheta,\eta}$ generates a $C_{0}$-semigroup, which is even analytic if and only if $\alpha_{1}/2\leq\beta_{1}\leq\alpha_{1},$ cf. [11]. On the other hand, $(J_{\vartheta,\eta}(t))_{t\geq 0}$ given by (4.9) is indeed a semigroup generated by $\mathcal{A}_{\vartheta,\eta}$ which follows by direct verification of the differential properties of $M,N$-functions in [29, p. 131] using the functional calculus. ∎

### v2

@@ Abstract
The coefficients of elastic and dissipative operators in a linear hyperbolic SPDE are jointly estimated using multiple spatially localised measurements. As the resolution level of the observations tends to zero, we establish the asymptotic normality of an augmented maximum likelihood estimator. The rate of convergence for the dissipative coefficients matches rates in related parabolic problems, whereas the rate for the elastic parameters also depends on the magnitude of the damping. The analysis of the observed Fisher information matrix relies upon the asymptotic behaviour of rescaled $M,N$-functions generalising the operator cosine and sine families appearing in the undamped wave equation. In contrast to the energetically stable undamped wave equation, the $M,N$-functions emerging within the covariance structure of the local measurements have additional smoothing properties similar to the heat kernel, and their asymptotic behaviour is analysed using functional calculus.
@@ 1 Introduction
for locations $x_{1},\dots,x_{N}\in\Lambda$ and $\gamma\in\{\alpha_{i},\beta_{j}|1\leq i\leq p;1\leq j\leq q\}$. The point spread functions [8, 9] $K_{\delta,x_{k}}$ are compactly supported functions taking non-zero values in an area centred around $x_{k}$ with radius $\delta$. Local measurements emerge naturally as they describe the physical limitation of measuring $u(t,x_{k})$, which, in general, is only possible up to a convolution with a point spread function. Local observations were introduced to the field of statistics for SPDEs in [5], where the authors investigated a stochastic heat equation with a spatially varying diffusivity. It was shown that the diffusivity at location $x_{k}\in\Lambda$ can be estimated based on a single local measurement process at $x_{k}$ as the resolution level $\delta$ tends to zero. The local observation scheme turned out to be robust under semilinearities [4, 3], multiplicative noise [19], discontinuities [32, 38] or lower-order perturbation terms [6, 37]. In contrast to the estimation of the diffusivity, the identifiability of transport or reaction coefficients necessarily requires an increasing amount $N\rightarrow\infty$ of measurements. In the recent contribution [41], the local measurement approach was extended to hyperbolic problems and the non-parametric wave speed in the undamped stochastic wave equation was estimated by relating the observed Fisher information to the energetic behaviour of an associated deterministic wave equation.
@@
with $N=N(\delta)$ measurements. The consistent estimation for $\vartheta_{i}$ holds if $N^{1/2}\delta^{-2\alpha_{i}+\alpha_{1}+\beta_{1}}\rightarrow\infty$, whereas $\eta_{j}$ can be estimated in the asymptotic regime $N^{1/2}\delta^{-2\beta_{j}+\beta_{1}}\rightarrow\infty$. In particular, estimating elastic coefficients is more difficult under higher dissipation, while damping coefficients are unaffected by the order of the elastic operator $A_{\vartheta}$ and their convergence rates reflect the rates obtained in advection-diffusion equations, cf. [6]. For the maximal number of non-overlapping observations $N\asymp\delta^{-d}$, our convergence rates match the rates obtained in the spectral approach up to specific boundary cases. In the weakly damped case, i.e. $\beta_{1}=0,$ we confirm the results in [24]. That is, the dependence of the time horizon $T$ of the asymptotic variance resembles the explosive, stable and ergodic cases of the maximum likelihood drift estimator for an Ornstein–Uhlenbeck process, cf. [22, Proposition 3.46]. In the structural damped case $(\beta_{1}>0)$, the asymptotic variance is of order $T^{-1}$ instead.
The central limited theorem follows conceptually through a martingale CLT and Slutsky’s lemma from the stochastic convergence of the rescaled observed Fisher information matrix. Verification of this convergence, however, is nontrivial and requires new approaches due to the structure of the semigroup associated with (1.1). It is characterised by $M,N$-functions, cf. [28, Chapter 1.7], which are generalisations of operator cosine and sin functions appearing in the solution to the stochastic wave equation. It is a-priori unclear whether the local measurement scheme is applicable in this setting since the involved $M,N$-functions inherit interacting properties from both the heat and the energetically stable wave equation. Smoothing and damping effects from the heat equation are introduced to the $M,N$-functions through the semigroup $(e^{tB_{\eta}})_{t\geq 0}$, while an oscillating behaviour is mirrored by terms of the form $(\cos(t(-A_{\vartheta}-B_{\eta}^{2}/4)^{1/2}))_{t\geq 0}$ and $(\sin(t(-A_{\vartheta}-B_{\eta}^{2}/4)^{1/2})(-A_{\vartheta}-B_{\eta}^{2}/4)^{-1/2})_{t\geq 0}$, respectively, which are the operator cosine and sine functions generated by $-A_{\vartheta}-B_{\eta}^{2}/4$. Consequently, both the elastic operator and the dissipative operator contribute to the oscillating component’s behaviour. The underlying solution operator combines phenomena from both heat and wave equations, which, surprisingly, results in a convergence proof that cannot directly combine the steps required in the asymptotic analysis for either of them. Instead, we exploit functional calculus for operators, which relies on a detailed analysis of the underlying $M,N$-functions and their rescaling when applied to localising kernel functions $K_{\delta,x}$. To the best of our knowledge, we are the first to utilise the theory of $M,N$-functions in a statistical field. Thus, we believe that our findings strengthen the existing toolbox of proof techniques and provide beneficial insights for future work in statistics of higher-order SPDEs.
@@ 2.2 The model
    Figure 2.1 displays a heatmap illustrating both the weakly and structurally damped plate equation in one spatial dimension. The solution of the SPDEs were approximated on a fine time-space grid using the finite difference scheme associated with the semi-implicit Euler-Maruyama method, see [25, Chapter 10]. Additional smoothing properties in the structurally damped case due to the dissipative operator $B_{\eta}$ result in an accelerated energetic decay in comparison to the weakly damped case.
@@ 2.3 Local measurements
A proper understanding of the covariance structure of the local measurements (2.5) is a key factor in the asymptotic analysis of the constructed estimator in Section 3. Denoting $L_{\vartheta,\eta}=-A_{\vartheta}-B_{\eta}^{2}/4$, the variance of a local measurement satisfies for instance
$\operatorname{Var}(u^{\Delta_{i}}_{\delta,x}(t))=\int_{0}^{t}\lVert e^{sB_{\eta}}\sin(sL_{\vartheta,\eta}^{1/2})L_{\vartheta,\eta}^{-1/2}(-\Delta)^{\alpha_{i}}K_{\delta,x}\rVert^{2}\mathop{}\!\mathrm{d}s,\quad 1\leq i\leq p,\quad 0<t\leq T.$ (2.6)
Intriguingly, phenomena from both heat and wave equations are infused in the representation of the variance (2.6). Indeed, in the heat equation (cf. [5]), the variance structure is determined by the semigroup $(e^{t\Delta})_{t\geq 0}$ generated by the Laplacian $\Delta$. It has a smoothing behaviour and induces the system’s energetic decay. In the wave equation (cf. [41]), the corresponding variance term relies on the oscillating operator sine function $(\sin(t(-\Delta)^{1/2})(-\Delta)^{-1/2})_{t\geq 0}$ generated by $-\Delta$. In (2.6), both effects are present through the differential operators $B_{\eta}$ and $L_{\vartheta,\eta}$ which leads to the study of $M,N$-functions generalising operator cosine and sine functions. When applied to localising functions $K_{\delta,x},$ a rescaling occurs, which takes both effects into account. Even though the limits coincide, asymptotic convergence of (2.6) for $\delta\rightarrow 0$ follows not through the integrated convergence of the heat kernel or the pointwise convergence (asymptotic equipartition of energy) known from the wave equation. Instead, we have to take a more subtle approach utilising functional calculus for operators as presented in Section 4.
@@ 3 The estimator
$\begin{pmatrix}\hat{\vartheta}_{\delta}\\ \hat{\eta}_{\delta}\end{pmatrix}=\begin{pmatrix}{\vartheta}\\ {\eta}\end{pmatrix}+\lVert K\rVert_{L^{2}(\operatorname{{\mathbb{R}}}^{d})}\mathcal{I}_{\delta}^{-1}\mathcal{M}_{\delta}$ (3.3)
@@
If the equation is weakly damped, i.e. if $\beta_{1}=0$ and $\eta_{1}<0$, then the term $-T\eta_{1}$ dominates the expression $(e^{T\eta_{1}}-T\eta_{1}-1)/2\eta_{1}^{2}$ in the asymptotic variance within (3.6) as $T\rightarrow\infty$. The converse is true in the amplified case with $\eta_{1}>0$. If $\eta_{1}=0$, the asymptotic variance of the augmented MLE for $\vartheta$ and $\eta_{1}$ depends on the time horizon through $T^{-2}$ as discussed in [41, Remark 5.8]. It mirrors the rate of convergence of the MLE in the ergodic, stable, and explosive case of the standard Ornstein–Uhlenbeck process as described in [22, Proposition 3.46].
@@
    Thus, the augmented MLE attains the convergence rate known from the spectral approach, see [24] if it is provided with the maximal number of spatial observations $N\asymp\delta^{-d}$. Interestingly, the limiting variance of $\hat{\eta}_{\delta}$ is independent of the kernel function $K$ similar to the augmented MLE for the first order transport coefficient in [6].
@@
    holds under the structural damping given in (3.8). The asymptotic variances between $\hat{\vartheta}_{\delta}$ and $\hat{\eta}_{\delta}$ coincide in the cases $\vartheta_{1}=\lVert\Delta_{0}K\rVert^{2}_{L^{2}(\operatorname{{\mathbb{R}}}^{d})}\lVert K\rVert_{L^{2}(\operatorname{{\mathbb{R}}}^{d})}^{-2}$ or $\vartheta_{1}=1$, respectively. The consistency and the varying convergence rates of the estimators are visualised in Figure 3.1. Based on the finite difference scheme within the semi-implicit Euler-Maruyama method [25, Chapter 10] (with $10000000\times 2000$ time-space grid points), we computed the root mean squared error (RMSE) for decreasing resolution level $\delta$ from 100 Monte Carlo runs, $N\asymp\delta^{-1}$ measurement locations and the kernel function $K(x)=\exp(-5/(1-x^{2}))\mathbf{1}(|x|<1)$. In the weakly damped case, it can be seen that the estimator for the elastic coefficient $\vartheta_{1}$ achieves a much quicker convergence rate than the estimator of the damping coefficient $\eta_{1}$. On the other hand, their rates are equal under structural damping. The asymptotic variances are attained in both cases.
@@
    hold. The convergence rates (3.9) are also obtained in the spectral regime, cf. [24, Theorem 1.1 and Theorem 4.1] under the consistency conditions (3.10) and (3.11). Given equality in these conditions, the authors further verified a logarithmic rate. We believe that the logarithmic rates in this boundary case are also valid in the local measurement approach given that less restrictive assumptions on the kernel $K$ are imposed, similar to [6, Proposition 5.4] in a related parabolic problem.
@@ 4 Proofs
The study of the operator $L_{\vartheta,\eta}$ and the convergence $L_{\vartheta,\eta,\delta,x}\rightarrow\bar{L}_{\vartheta,\eta}$ is of particular importance as $L_{\vartheta,\eta}$ appears in the solution to the coupled second-order system (2.1) through the $M,N$-functions (generalisations of operator cosine and sine functions) introduced in Lemma 4.4 below. It captures the interaction between the elasticity operator $A_{\vartheta}$ and the dissipative operator $B_{\eta}.$ The asymptotically dominating term of $L_{\vartheta,\eta,\delta,x}$ determines the limiting object $\bar{L}_{\vartheta,\eta}$. We will thus start by some fundamental observations regarding the differential operator $L_{\vartheta,\eta}$ which plays a crucial role in the following Sections 4.1 and 4.2.
We will frequently use that $L_{\vartheta,\eta}$ is an (unbounded) normal operator with spectrum $\sigma(L_{\vartheta,\eta})$ and the resolution of identity $E$ (cf. [33, Chapter 13]). By the functional calculus for normal unbounded operators, we can define the operator $f(L_{\vartheta,\eta})\coloneqq\int_{\sigma(L_{\vartheta,\eta})}f(\lambda)\mathop{}\!\mathrm{d}E(\lambda)$ on the domain
@@
Suppose $z\in\dot{H}^{2\alpha_{1}}(\Lambda_{\delta,x})$ such that $z_{\delta,x}\in\dot{H}^{2\alpha_{1}}(\Lambda)\subset\mathcal{D}_{f}$. Then, the claim follows immediately for $f(x)=x$ by differentiating $z_{\delta,x}$ and from the definition of $L_{\vartheta,\eta,\delta,x}$, see also [4, Lemma 16]. Using (i) and (iii) of [34, Proposition 5.15], the result can be extended to measurable $f:\mathbb{C}\rightarrow\mathbb{C}\cup\{\pm\infty\}$ by first passing to the associated resolution of the identities of $L_{\vartheta,\eta},B_{\eta},A_{\vartheta}$ and $L_{\vartheta,\eta,\delta,x},B_{\eta,\delta,x},A_{\vartheta,\delta,x}$ respectively, and interpreting the localisation as a bounded linear operator from $L^{2}(\Lambda_{\delta,x})$ to $L^{2}(\Lambda)$. ∎
@@
If $L_{\vartheta,\eta}$ is not a positive operator and has non-positive eigenvalues, any choice of the operator root is a complex-valued operator. This is particularly the case if the structural damping exceeds the threshold $\alpha_{1}=2\beta_{1}$. Consequently, the associated family of $M,N$-functions in the following subsection is again complex-valued. Thus, inner products hereinafter are associated with complex Hilbert spaces. However, this does not influence the asymptotic results of Section 4.2 due to the convergence $L_{\vartheta,\eta,\delta,x}\rightarrow\bar{L}_{\vartheta,\eta}$ to a positive limiting operator $\bar{L}_{\vartheta,\eta}$.
@@ 4.1 Properties of generalised cosine and sine operator functions
Note that by [34, Theorem 5.9] all of the appearing operators $e^{B_{\eta}t/2}$, $\cos(tL_{\vartheta,\eta}^{1/2})$, $\sin(tL_{\vartheta,\eta}^{1/2})$, $B_{\eta}$ and $L_{\vartheta,\eta}^{-1/2}$ in (4.2) and (4.3) are well-defined and even commute on the smallest occurring domain as they are all based on the same underlying Laplace operator. By direct computation, one can now verify that the conditions (M1)-(M4) in [28, Definition 1.7.2] are satisfied by $M(t)$ and $N(t)$ from (4.2) and (4.3) using the functional calculus. ∎
@@
The unique positive self-adjoint operator root of the positive self-adjoint operator $L_{\vartheta,\eta}$ is well-defined and exists by [34, Proposition 5.13]. Thus, in view of Lemma 4.4, the $M,N$-functions can each be interpreted as the applications of a real-valued function to the underlying Laplace operator on a bounded spatial domain. In particular, by [34, Theorem 5.9] the $M,N$-functions are self-adjoint. ∎
@@
Let $\beta_{1}>0.$ We only prove (4.7), since the argument for (4.6) is similar, using additionally that $L_{\vartheta,\eta,\delta,x}$ commutes with $(-\Delta_{\delta,x})^{\gamma}$, $\operatorname{ord}(L_{\vartheta,\eta,\delta,x})=2\alpha_{1}$ and a bound of $L_{\vartheta,\eta,\delta,x}$ in terms of its leading term $(-\Delta_{\delta,x})^{\alpha_{1}}.$ Let $(e_{k})_{k\in\operatorname{{\mathbb{N}}}}$ form an orthonormal basis of $(-\Delta)$ in $L^{2}(\Lambda)$ with eigenvalues $\lambda_{k}>0$. Then, there exists a constant $c(\Lambda)$ such that $\lambda_{k}\geq c(\Lambda)$ for all $k\geq 1$, see [35, Proposition 5.2 and Corollary 5.3]. We consider the most involved case, that is, $\eta_{1}<0$ and $\eta_{2},\dots,\eta_{q}>0$. Consequently, $B_{\eta}$ will, in general, not be a negative operator, but there exists $y_{0}>0$ such that for all $y\geq y_{0}$, we have
@@
Thus, all eigenvalues of the operator difference $B_{\eta,\delta,x}-\delta^{2\beta_{1}}C_{1}$ are negative and the difference is bounded by the differential operator $\eta_{1}(-\Delta_{\delta,x})^{\beta_{1}}/2$ in the sense that
@@
The theory of cosine operator functions was developed by [36] and led to general deterministic solution theory for undamped second-order abstract Cauchy problems. By substituting the time derivative as its own variable, it is possible to rewrite a second-order abstract Cauchy problem as a first-order abstract Cauchy problem in two components. The associated strongly continuous semigroup then lives on a product of Hilbert spaces called the phase-space; see [7, Chapter 3.14], [29] and [28, Chapter 0.3]. The same remains true under suitable assumptions on the elastic and dissipation operator within a damped abstract second-order Cauchy problem [28, Chapter 1.7].
@@
It is well-known that in the special case where both $A_{\vartheta}$ and $B_{\eta}$ are strictly negative operators $\mathcal{A}_{\vartheta,\eta}$ generates a $C_{0}$-semigroup, which is even analytic if and only if $\alpha_{1}/2\leq\beta_{1}\leq\alpha_{1},$ cf. [11]. On the other hand, $(J_{\vartheta,\eta}(t))_{t\geq 0}$ given by (4.9) is indeed a semigroup generated by $\mathcal{A}_{\vartheta,\eta}$ which follows by direct verification of the differential properties of $M,N$-functions in [28, p. 131] using the functional calculus. ∎